# Splitting obstructions and $\mathbb{Z}_{2}$ invariants in time-reversal symmetric topological insulators

### v1

@@ 2.1. Continuous parallel transport
In this subsection, we investigate the problem of constructing a unitary-valued map $U:\mathbb{T}^{2}\to\mathcal{U}(\mathcal{H})$ associated with a given projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$, such that
$P(k)=U(k)\,P(0)\,U(k)^{-1}.$
@@
A standard approach to this problem relies on parallel transport (see, e.g., [Cornean_2017, monaco2023az]), which requires at least $C^{1}$ regularity of the map $k\mapsto P(k)$. Here, we present a method based on the Kato–Nagy intertwiner, which applies already in the merely continuous setting and thus provides a genuine generalization of parallel transport to lower regularity.
@@
If $P,Q\in\operatorname{Proj}_{n}(\mathcal{H})$ are two projections with $\|P-Q\|<1$, then
@@
This formula will be used extensively; its proof can be found in [kato2013perturbation].
Proposition 2.2 (Kato–Nagy 1-d construction: continuous parallel transport).
If $P:\mathbb{T}^{1}\to\operatorname{Proj}_{m}(\mathcal{H})$ is a continuous projection-valued map, then there exists a continuous unitary-valued map $\tilde{U}:\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$ such that
$P(k)=\tilde{U}(k)P(0)\tilde{U}(k)^{-1}.$
If $P$ is time-reversal symmetric, it is possible to choose $\tilde{U}$ so that $T\tilde{U}(k)=\tilde{U}(-k)T$.
@@
Divide $0,\pi]$ into a finite number of intervals $I_{j}=[t_{j},t_{j+1}]$ with $t_{0}=0$, $t_{J}=\pi$, such that $\|P(t_{j})-P(t)\|<1$ for all $t\in[t_{j},t_{j+1}]$. For $t\in I_{j}$, define $U_{t_{j}}^{t}=U_{P(t_{j})\to P(t)}$ according to Theorem 2.1. The unitary equivalence between $P(0)$ and $P(t)$ is then
$U(t)=U_{t_{j}}^{t}U_{t_{j-1}}^{t_{j}}\cdots U_{t_{0}}^{t_{1}}.$
If $P$ is time-reversal symmetric, extend $U$ to $\pi,0]$ via U(t)=TU(t)TU(t)=TU(-t)T; otherwise, replicate the same procedure for $t\in[-\pi,0]$. In both cases, $U:[-\pi,\pi]\to\mathcal{U}(\mathcal{H})$ satisfies $P(t)=U(t)P(0)U(t)^{-1}$ for all $t$.
This $U$ is generally not periodic in $t$, so consider $U(-\pi)^{-1}U(\pi)$, which commutes with $P(0)$. If $P$ is symmetric, it also satisfies $TU(-\pi)^{-1}U(\pi)=\left[U(-\pi)^{-1}U(\pi)\right]^{-1}T$. By the spectral theorem, there exists a self-adjoint operator $L$ commuting with $P(0)$ such that $U(-\pi)^{-1}U(\pi)=e^{iL}$. If $P$ is symmetric, $L$ can also be chosen to commute with $T$. Finally, define
$\tilde{U}(t)=U(t)e^{-itL/2\pi}.$
This construction ensures that $\tilde{U}$ is periodic, continuous, intertwines $P(0)$ with $P(t)$, and respects the symmetry constraint when $P$ does. ∎
Proposition 2.3 (Kato–Nagy 2-d construction).
If $P:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$ is a continuous projection-valued map, then there exists a continuous unitary-valued map $U:[-\pi,\pi]\times\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$, periodic in the second component but not in the first, such that
$P(t,k_{2})=U(t,k_{2})P(0,0)U(t,k_{2})^{-1}.$
@@
Divide $0,\pi]$ into a finite number of intervals $t_{j},t_{j+1}]$ with $t_{0}=0$, $t_{J}=2\pi$, such that $\|P(t_{j},k_{2})-P(t,k_{2})\|<1/2$ for all $(t,k_{2})\in[t_{j},t_{j+1}]\times\mathbb{T}^{1}$. Define $U_{t_{j}}^{t}(k_{2})=U_{P(t_{j},k_{2})\to P(t,k_{2})}$ via Theorem 2.1.
To define the unitary equivalence between $P(0,0)$ and $P(t,k_{2})$, first apply Proposition 2.2 to the 1-d projection-valued map $P(0,k_{2})$ to construct a periodic (and possibly symmetric) unitary-valued map $\tilde{U}(k_{2})$ with $P(0,k_{2})=\tilde{U}(k_{2})P(0,0)\tilde{U}(k_{2})^{-1}$. Then, for $t\in[t_{j},t_{j+1}]$, set
$U(t,k_{2})=U_{t_{j}}^{t}(k_{2})U_{t_{j-1}}^{t_{j}}(k_{2})\cdots U_{t_{0}}^{t_{1}}(k_{2})\tilde{U}(k_{2}).$
@@ 2.2. Chern number through the matching matrices
In the following sections, it will be essential to adopt an alternative formulation of the Chern number: compare [Cornean_2017].
@@
Let $P:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$ be a continuous, not necessarily symmetric, projection-valued map, and let $U:[-\pi,\pi]\times\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$ be a continuous unitary-valued map, periodic in the second argument, such that
$P(t,k_{2})=U(t,k_{2})P(0,0)U(t,k_{2})^{-1}.$
Then $U(-\pi,k_{2})^{-1}U(\pi,k_{2})$ commutes with $P(0,0)$, so after choosing an orthonormal basis $\{v_{j}\}_{j=1,\dots,n}$ spanning $\operatorname{Ran}(P(0,0))$, we define unitary matrices $\alpha(k_{2})\in M_{n,n}(\mathbb{C})$, called matching matrices, depending periodically and continuously on $k_{2}$, via
$\alpha_{a,b}(k_{2})=\langle v_{a},U(-\pi,k_{2})^{-1}U(\pi,k_{2})v_{b}\rangle.$
The Chern number is then the integer obtained as the winding number of $\det(\alpha)$ :
$\operatorname{Ch}(P)=[\det(\alpha(\cdot))].$
If $\dim(\operatorname{Ran}(P(k)))=\infty$, we set $\operatorname{Ch}(P)=0$.
In the remainder of this section, we derive from the previous definition some fundamental properties of the Chern number. Although these results are well known, we believe that presenting their proofs in the notation introduced above may help the reader become familiar with the techniques employed in the following sections.
@@
The above definition is independent of the choice of the unitary operator $U$. Moreover, if two projection-valued maps are unitarily equivalent or homotopic, their Chern numbers coincide.
@@
Well-posedness.  First, we show that the Chern number does not depend on the choice of the orthonormal basis $\{v_{j}\}$. If $\{v_{j}^{\prime}\}$ is another basis, it induces a different $\alpha^{\prime}(k_{2})$, which is unitarily equivalent to $\alpha(k_{2})$ via $A_{a,b}=\langle v_{a},v_{b}^{\prime}\rangle$, so
@@
continuity in $t$ and periodicity in $k_{2}$ imply that $\det\beta(t,\cdot)]$ is constant in $t$. By Corollary A.5 and $\det(AB)=\det(A)\det(B)$, we obtain
@@
Unitary invariance.  Let $V(k)$ implement a unitary equivalence between two projection-valued maps $P_{0}(k)$ and $P_{1}(k)$:
$P_{1}(k)=V(k)P_{0}(k)V(k)^{-1}.$
Given a Kato-Nagy unitary $U$ such that
@@
$\alpha_{1}]_{a,b}(k_{2})=\langle v_{a}^{1},V(0,0)U(-\pi,k_{2})^{-1}V(-\pi,k_{2})^{-1}V(\pi,k_{2})U(\pi,k_{2})V(0,0)^{-1}v_{b}^{1}\rangle,$
periodicity of $V$ in $k_{2}$ and the relation $v_{j}^{0}:=V(0,0)^{-1}v_{j}^{1}$ lead to
$\alpha_{1}]_{a,b}(k_{2})=\langle v_{a}^{0},U(-\pi,k_{2})^{-1}U(\pi,k_{2})v_{b}^{0}\rangle.$
@@
Homotopy invariance.  Consider a homotopy $P_{t}(k)$ of projection-valued maps. By compactness of $0,1]\times\mathbb{T}^{2}$ and continuity, we can partition $0,1]=\bigcup_{j=0}^{J-1}[t_{j},t_{j+1}]$ such that
@@
Then, for each $j$, the Kato-Nagy unitary
$U_{j}(k)=\big[P_{t_{j+1}}(k)P_{t_{j}}(k)+(\mathds{1}-P_{t_{j+1}}(k))(\mathds{1}-P_{t_{j}}(k))\big]\big[\mathds{1}-(P_{t_{j}}(k)-P_{t_{j+1}}(k))^{2}\big]^{-1/2}$
implements a unitary equivalence between $P_{t_{j}}$ and $P_{t_{j+1}}$. By the previous item, Chern numbers are preserved along each step, yielding
@@
The Chern number is additive: if $P:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$ and $Q:\mathbb{T}^{2}\to\operatorname{Proj}_{m}(\mathcal{H})$ satisfy $PQ\equiv 0$, then for $P+Q:\mathbb{T}^{2}\to\operatorname{Proj}_{n+m}(\mathcal{H})$,
@@
Then we extend the operator $X:[0,2\pi]\times\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$ by linearity with
$X(u)=\begin{cases}U(u)\quad\mbox{if}\quad Pu=u\\ V(u)\quad\mbox{if}\quad Qu=u\\ Y(u)\quad\mbox{if}\quad Pu=Qu=0\end{cases}$
to obtain a unitary-valued map with
@@
Now we compute the Chern number of $P+Q$ using a basis $\{v_{j}\}_{j\in\{1,\cdots,m+n\}}$ where the first $n$ span $P(0,0)$ and the other span $Q(0,0)$. In this way we have
@@
where $\beta:\mathbb{T}^{2}\to U(n)$ and $\gamma:\mathbb{T}^{2}\to U(m)$ are such that
$\beta(k_{2})]_{a,b}=\left\langle v_{a},U(-\pi,k_{2})^{-1}U(\pi,k_{2})v_{b}\right\rangle\quad\text{and}\quad[\gamma(k_{2})]_{a,b}=\left\langle v_{n+a},V(-\pi,k_{2})^{-1}V(\pi,k_{2})v_{n+b}\right\rangle\,.$
So using Corollary A.5 we obtain the thesis:
$\operatorname{Ch}(P+Q)=[\det(\alpha(k_{2}))]=[\det(\beta(k_{2}))\det(\gamma(k_{2}))]=[\det(\beta(k_{2}))]+[\det(\gamma(k_{2}))].\qed$
3. The topological invariant
In a milestone paper [kane2005z], Kane and Mele demonstrated that it is possible to associate a $\mathbb{Z}_{2}$-valued topological invariant to a specific discrete honeycomb model. The methods introduced in that work were subsequently extended to a broader class of models in several later papers, including [fu2007topological], [freed2013twisted], [graf2013bulk], [de2015classification], [fiorenza2016z], and [monaco2015symmetry], each providing a distinct formulation of the invariant and a different interpretation of the physical information it encodes. In this section, we will present yet another formulation of the invariant, starting from a continuous projection-valued map
$P:\mathbb{T}^{2}\to\operatorname{Proj}_{m}(\mathcal{H})$
subject to the symmetry constraint described in Definition 1.4. We also note, as Lemma A.1 guarantees, that $\dim(\operatorname{Ran}(P(0,0)))=m$ is always even; hence, from this point onward, we set $m=2n$.

### v2

@@ 2. Continuous parallel transport
One typical approach to the construction of a frame for a given projection-valued map $\{P(k)\}_{k\in\mathbb{R}^{2}}$ is based on the following observation: if one can unitarily intertwine two projections in the family, i.e. if $P(k^{\prime})=U\,P(k)\,U^{-1}$ for some $k,k^{\prime}\in\mathbb{R}^{2}$, then an orthonormal basis for $\operatorname{Ran}P(k)$ can be “transported”, via the unitary $U$, to yield a basis for $\operatorname{Ran}P(k^{\prime})$. Therefore, in this Section we investigate the problem of constructing a unitary-valued map $U:\mathbb{T}^{2}\to\mathcal{U}(\mathcal{H})$ associated with a given projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$, such that
(5) $P(k)=U(k)\,P(0)\,U(k)^{-1}.$
If $P$ is further time-reversal symmetric, we will also require that
(6) $TU(k)=U(-k)T,\quad\forall k\in\mathbb{T}^{2},$
so that the time-reversal symmetry constraint can be preserved also at the level of symmetries.
As we shall see, it is in general impossible to find such a map $U$ that is continuous and periodic in both variables; periodicity can be achieved only in one of them, yielding a continuous map defined on $\pi,\pi]\,\times\mathbb{T}^{1}$ which satisfies (5) and, in the presence of time-reversal symmetry, also (6).
@@
A standard approach to this problem relies on parallel transport, which is widely used in the literature on symmetric projection-valued maps (see, e.g., [cornean2016construction, cances2017robust, Cornean_2017, cornean2019parseval, gontier2019numerical, gontier2022symmetric, monaco2023az]). This method requires at least $C^{1}$ regularity of the map $k\mapsto P(k)$. Here, also in view of Remark 1.10, we present a method based on the Kato–Nagy intertwiner, which applies already in the merely continuous setting and thus provides a genuine generalization of parallel transport to lower regularity, in the context of continuous projection-valued maps on the torus $\mathbb{T}^{d}$. The whole construction is based on the following Theorem, a proof of which can be found in [kato2013perturbation]:
@@
If $P,Q\in\operatorname{Proj}_{r}(\mathcal{H})$ are two projections with $\|P-Q\|<1$, then
@@
The next result shows the existence of a periodic and time-reversal symmetric intertwiner for $1$-dimensional periodic restrictions of the projection-valued map.
Proposition 2.2.
If $P:\mathbb{T}^{1}\to\operatorname{Proj}_{m}(\mathcal{H})$ is a continuous projection-valued map, then there exists a continuous unitary-valued map $U:\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$ such that
$P(k)=U(k)P(0)U(k)^{-1}.$
If $P$ is time-reversal symmetric, it is possible to choose $U$ so that TU(k)=U(k)TTU(k)=U(-k)T.
@@
Divide $0,\pi]$ into a finite number of intervals $I_{j}=[t_{j},t_{j+1}]$ with $t_{0}=0$, $t_{J}=\pi$, such that $\|P(t_{j})-P(t)\|<1$ for all $t\in[t_{j},t_{j+1}]$. For $t\in I_{j}$, define $U_{t_{j}}^{t}=U_{P(t_{j})\to P(t)}$ according to Theorem 2.1. A unitary equivalence between $P(0)$ and $P(t)$, $t\in[0,\pi]$ is obtained via
$\tilde{U}(t):=U_{t_{j}}^{t}U_{t_{j-1}}^{t_{j}}\cdots U_{t_{0}}^{t_{1}}.$
If $P$ is time-reversal symmetric, extend $\tilde{U}$ to $\pi,0]$ via $\tilde{U}(t)=T\tilde{U}(-t)T$; otherwise, replicate the same procedure for $t\in[-\pi,0]$. In both cases, $\tilde{U}:[-\pi,\pi]\to\mathcal{U}(\mathcal{H})$ satisfies $P(t)=\tilde{U}(t)P(0)\tilde{U}(t)^{-1}$ for all $t\in[-\pi,\pi]$.
This $\tilde{U}$ is in general not periodic in $t$, so consider $\tilde{U}(-\pi)^{-1}\tilde{U}(\pi)$, which commutes with $P(0)$. If $P$ is symmetric, it also satisfies $T\left[\tilde{U}(-\pi)^{-1}\tilde{U}(\pi)\right]=\left[\tilde{U}(-\pi)^{-1}\tilde{U}(\pi)\right]^{-1}T$. By the spectral theorem, there exists a self-adjoint operator $L$ commuting with $P(0)$ such that $\tilde{U}(-\pi)^{-1}\tilde{U}(\pi)=e^{iL}$. If $P$ is symmetric, $L$ can also be chosen to commute with $T$. Finally, define
$U(t)=\tilde{U}(t)e^{-itL/2\pi}.$
This construction ensures that $U(\pi)=U(-\pi)$, and therefore $U$ can be continuously extended to the whole $\mathbb{R}$ in a periodic way, maintaining the intertwining condition between $P(0)$ and $P(t)$, and respecting the symmetry constraint when $P$ does. ∎
As mentioned above, the latter construction fails to produce a fully-periodic unitary intertwiner, if one moves to $d=2$.
Proposition 2.3.
If $P:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$ is a continuous projection-valued map, then there exists a continuous unitary-valued map $U:[-\pi,\pi]\times\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$, periodic in the second variable, such that
$P(t,k_{2})=U(t,k_{2})P(0,0)U(t,k_{2})^{-1}\quad\forall\;(t,k_{2})\in[-\pi,\pi]\times\mathbb{T}^{1}.$
@@
Divide $0,\pi]$ into a finite number of intervals $t_{j},t_{j+1}]$ with $t_{0}=0$, $t_{J}=2\pi$, such that $\max_{k_{2}\in\mathbb{T}^{1}}\|P(t_{j},k_{2})-P(t,k_{2})\|<1/2$ for all $(t,k_{2})\in[t_{j},t_{j+1}]\times\mathbb{T}^{1}$. Define $U_{t_{j}}^{t}(k_{2})=U_{P(t_{j},k_{2})\to P(t,k_{2})}$ via Theorem 2.1.
To define the unitary equivalence between $P(0,0)$ and $P(t,k_{2})$, first apply Proposition 2.2 to the 1-d projection-valued map $P(0,k_{2})$ to construct a periodic (and possibly symmetric) unitary-valued map $V\colon\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$ with $P(0,k_{2})=V(k_{2})P(0,0)V(k_{2})^{-1}$. Then, for $t\in[t_{j},t_{j+1}]$, set
$U(t,k_{2}):=U_{t_{j}}^{t}(k_{2})U_{t_{j-1}}^{t_{j}}(k_{2})\cdots U_{t_{0}}^{t_{1}}(k_{2})V(k_{2}).$
@@
3. Review: frames in absence of time-reversal symmetry
In this Section, we momentarily forgo the constraint of time-reversal symmetry, and review how the construction of continuous and periodic frames for 2-dimensional periodic projection-valued maps gives rise to an integer valued topological obstruction, namely the Chern number of the family of projections. We follow the method presented e.g. in [cornean2019parseval, monaco2023topology]: we refer the reader there and to the references therein for a more detailed account. The present presentation has the advantage to fully remain in the continuous regularity class, and to set up notions and notation for the generalization to the time-reversal symmetric setting.
The aim of this Section is thus to introduce two constructions that will play a central role throughout the text, related to a rank-$m$ projection-valued map $\{P(k)\}_{k\in\mathbb{T}^{2}}$. The first provides an equivalent formulation of the Chern number in terms of the matching matrices defined from the intertwining unitaries constructed in Proposition 2.3. The second provides a way to build a continuous frame $\{v_{j}(t,k_{2})\}_{1\leq j\leq m}$ for $P$, $(t,k_{2})\in[-\pi,\pi]\times\mathbb{T}^{1}$, that is “as periodic as possible” – precisely, pseudo-periodic, in a sense akin to that introduced in Definition 1.6.
Definition 3.1 (Pseudo–periodic frame).
Let $P$ be a rank-$m$ projection-valued map in the sense of Definition 1.4. A frame $\{v_{j}\}_{1\leq j\leq m}$ is called pseudo-periodic if the vectors define continuous maps from $\pi,\pi]\times\mathbb{T}^{1}$ to $\mathcal{H}$ that satisfy
  1. (1)
    periodicity on m1m-1 elements, i.e.
    $v_{j}(\pi,k_{2})=v_{j}(-\pi,k_{2})\quad\forall k_{2}\in\mathbb{T}^{1},\qquad\forall j\neq 1;$
  2. (2)
    periodicity up to a phase on $1$ element, , i.e. for some $h\in\mathbb{Z}$
    $v_{1}(\pi,k_{2})=e^{ihk_{2}}v_{1}(-\pi,k_{2})\qquad\forall k_{2}\in\mathbb{T}^{1}.$
Moreover, we will relate the lack of periodicity of the vector $v_{1}$ as above to the Chern number: $h=\operatorname{Ch}(P)\in\mathbb{Z}$.
@@ 3.1. Chern number through the matching matrices
Let $P:\mathbb{T}^{2}\to\operatorname{Proj}_{m}(\mathcal{H})$ be a continuous, not necessarily symmetric, projection-valued map, and let $U:[-\pi,\pi]\times\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$ be a continuous unitary-valued map as in Proposition 2.3, that is, periodic in the second argument and such that
$P(t,k_{2})=U(t,k_{2})P(0,0)U(t,k_{2})^{-1}\quad\forall(t,k_{2})\in[-\pi,\pi]\times\mathbb{T}^{1}.$
Then $U(-\pi,k_{2})^{-1}U(\pi,k_{2})$ commutes with $P(0,0)$, so after choosing an orthonormal basis $\{v_{j}\}_{1\leq j\leq m}$ spanning $\operatorname{Ran}P(0,0)$ we can define
$\alpha(k_{2})_{a,b}:=\langle v_{a},U(-\pi,k_{2})^{-1}U(\pi,k_{2})v_{b}\rangle\,,\quad a,b\in\{1,\ldots,m\}.$
The matrices $\alpha(k_{2})=[\alpha(k_{2})_{a,b}]_{1\leq a,b\leq m}\in M_{m,m}(\mathbb{C})$, called matching matrices, are then unitary, and depend periodically and continuously on $k_{2}\in\mathbb{T}^{1}$.
@@
If $m\in\mathbb{N}$, the Chern number of the rank-$m$ projection-valued map $P$ is the integer $\operatorname{Ch}(P)\in\mathbb{Z}$ obtained as the winding number of $\det\alpha\colon\mathbb{T}^{1}\to\mathrm{U}(1)$ (compare Definition A.2):
$\operatorname{Ch}(P):=[\det(\alpha(\cdot))].$
If $m=\dim\operatorname{Ran}P(k)=\infty$, we set $\operatorname{Ch}(P):=0$.
In the remainder of this Section, we derive from the previous definition some fundamental properties of the Chern number.
@@
The above definition is independent of the choice of the unitary operator $U$ and of the basis $\{v_{j}\}$ for $\operatorname{Ran}P(0,0)$. Moreover, if two projection-valued maps are Murray–von Neumann equivalent, unitarily equivalent or homotopic, their Chern numbers coincide.
@@
Well-posedness.  First, we show that the Chern number does not depend on the choice of the orthonormal basis $\{v_{j}\}$. If $\{v_{j}^{\prime}\}$ is another basis, it induces a different $\alpha^{\prime}(k_{2})$, which is unitarily equivalent to $\alpha(k_{2})$ via $A_{a,b}:=\langle v_{a},v_{b}^{\prime}\rangle$; consequently
@@
continuity in $t$ and periodicity in $k_{2}$ imply that $\det\beta(t,\cdot)]$ is constant in $t$. By Proposition A.5 and $\det(AB)=\det(A)\det(B)$, we obtain
@@
Murray–von Neumann and unitary invariance.  Let $V(k)$ implement a Murray–von Neumann equivalence or a unitary equivalence between two projection-valued maps $P_{0}(k)$ and $P_{1}(k)$. In both cases we have
$P_{1}(k_{1},k_{2})=V(k_{1},k_{2})P(k_{1},k_{2})V(k_{1},k_{2})^{*}$
From proposition 2.3, there exists a unitary–valued map $U$, periodic in the second variable, such that
@@
$\alpha_{1}(k_{2})_{a,b}=\langle v_{a}^{1},V(0,0)U(-\pi,k_{2})^{-1}V(-\pi,k_{2})^{*}V(\pi,k_{2})U(\pi,k_{2})V(0,0)^{*}v_{b}^{1}\rangle,$
periodicity of $V$ in $k_{2}$ and the relation $v_{j}^{0}:=V(0,0)^{*}v_{j}^{1}$ lead to
$\alpha_{1}(k_{2})_{a,b}=\langle v_{a}^{0},U(-\pi,k_{2})^{-1}U(\pi,k_{2})v_{b}^{0}\rangle.$
@@
Homotopy invariance.  Consider a homotopy $P_{t}(k)$ of projection-valued maps. By compactness of $0,1]\times\mathbb{T}^{2}$ and continuity, we can partition $0,1]=\bigcup_{j=0}^{J-1}[t_{j},t_{j+1}]$ in a way such that
@@
Then, for each $j$, the Kato-Nagy unitary $U_{j}(k):=U_{P_{t_{j}}(k)\to P_{t_{j+1}}(k)}$ implements a unitary equivalence between $P_{t_{j}}$ and $P_{t_{j+1}}$. By the previous item, Chern numbers are preserved along each step, yielding
@@
The Chern number is additive: if $P:\mathbb{T}^{2}\to\operatorname{Proj}_{r}(\mathcal{H})$ and $Q:\mathbb{T}^{2}\to\operatorname{Proj}_{m}(\mathcal{H})$ satisfy $PQ\equiv 0$, then $P+Q:\mathbb{T}^{2}\to\operatorname{Proj}_{r+m}(\mathcal{H})$ is still a continuous periodic projection-valued map satisfying
@@
We set
$X(u):=\begin{cases}U(u)\quad\mbox{if}\quad Pu=u\\ V(u)\quad\mbox{if}\quad Qu=u\\ Y(u)\quad\mbox{if}\quad Pu=Qu=0\end{cases}$
and extend the definition to $X:[0,2\pi]\times\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$ by linearity, to obtain a unitary-valued map with
@@
Now we compute the Chern number of $P+Q$ using a basis $\{v_{j}\}_{j\in\{1,\cdots,r+m\}}$ where the first $r$ vectors span $P(0,0)$ and the other $m$ span $Q(0,0)$. In this way we have
@@
where $\beta:\mathbb{T}^{2}\to\mathrm{U}(r)$ and $\gamma:\mathbb{T}^{2}\to\mathrm{U}(m)$ are such that
$\beta(k_{2})_{a,b}=\left\langle v_{a},U(-\pi,k_{2})^{-1}U(\pi,k_{2})v_{b}\right\rangle\quad\text{and}\quad\gamma(k_{2})_{a,b}=\left\langle v_{r+a},V(-\pi,k_{2})^{-1}V(\pi,k_{2})v_{r+b}\right\rangle\,.$
Applying Proposition A.5 we obtain the desired equality:
$\operatorname{Ch}(P+Q)=[\det\alpha]=[\det\beta\cdot\det\gamma]=[\det\beta]+[\det\gamma]=\operatorname{Ch}(P)+\operatorname{Ch}(Q).\qed$
3.2. Pseudo-periodic frames
The next statement exhibits the construction of a pseudo-periodic frame for a continuous periodic projection-valued map, and expresses the Chern number as the obstruction to finding a fully periodic frame.
Theorem 3.5 (Existence of a pseudo-periodic frame).
Consider a projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{m}(\mathcal{H})$. Then it is always possible to construct a pseudo-periodic frame $\{v_{j}(k)\}_{1\leq j\leq m}$ for $P$ as in Definition 3.1, with $h=\operatorname{Ch}(P)\in\mathbb{Z}$.
Proof.
If $\{v_{j}\}_{1\leq j\leq m}$ is an orthonormal basis of $\operatorname{Ran}P(0,0)$, then $\{u_{j}(t,k_{2})=U(t,k_{2})v_{j}\}_{1\leq j\leq m}$ constitutes an orthonormal basis of $P(t,k_{2})$ that depends continuously in $t$ and $k_{2}$. This is periodic in $k_{2}$ but in general will not be periodic in $t$. The aperiodicity is once again measured by the matching matrix $\alpha(k_{2})\in\mathrm{U}(m)$, with:
$u_{j}(\pi,k_{2})=\sum_{a=1}^{m}[\alpha(k_{2})]_{j,a}u_{j}(-\pi,k_{2}).$
Since $\det\alpha]=\operatorname{Ch}(P)$ the matrix:
$\beta(k_{2})=\begin{pmatrix}e^{i\operatorname{Ch}(P)k_{2}}&0\\ 0&\mathds{1}_{n-1}\end{pmatrix}\alpha(k_{2})^{-1}$
is such that $\det\beta]=0$. By Theorem A.6, there exists a homotopy $\beta_{s}(k_{2})$ with $\beta_{1}(k_{2})=\beta(k_{2})$ and $\beta_{0}(k_{2})=\mathds{1}$. Therefore, the vectors
$v_{j}(t,k_{2})=\sum_{a=1}^{m}\left[\beta_{(t+\pi)/2\pi}(k_{2})\right]_{j,a}u_{j}(t,k_{2})\quad\mbox{for}\quad j\in\{1,\cdots,n\}$
constitute an orthonormal basis of $\operatorname{Ran}P(t,k_{2})$ and are such that:
$v_{j}(\pi,k_{2})$ $=\sum_{a=1}^{m}[\beta(k_{2})]_{j,a}u_{a}(\pi,k_{2})=\sum_{a,b=1}^{m}[\beta(k_{2})]_{j,a}\alpha_{a,b}(k_{2})u_{b}(-\pi,k_{2})$
$=\sum_{a=1}^{m}[\beta(k_{2})\alpha(k_{2})]_{j,a}u_{a}(-\pi,k_{2})=\begin{cases}e^{i\operatorname{Ch}(P)k_{2}}v_{1}(-\pi,k_{2})&\mbox{for }j=1\,,\\ v_{j}(-\pi,k_{2})&\mbox{otherwise}.\end{cases}\qed$
4. From the symmetric matching matrices to the $\mathbb{Z}_{2}$-valued invariant
Analogously to what happens in the absence of time-reversal symmetry, one can associate to a time-reversal symmetric projection-valued map on the torus a topological invariant, with values in $\mathbb{Z}_{2}$, obtained from its matching matrices. In particular, the fact that the projection-valued map is subject to time-reversal symmetry induces a corresponding symmetry on the associated matching matrices, provided these are computed from a pseudo-periodic and symmetric frame. This additional symmetry and its consequences are discussed in the first subsection, which essentially reviews some definitions and results contained in [Cornean_2017]. The second subsection defines the $\mathbb{Z}_{2}$ topological invariant, showing that it truly depends only on the family of projections and not on the choices made to define it from the matching matrices. Our definition can be related to the one proposed by Graf and Porta in [graf2013bulk], as discussed in Remark 4.5, and therefore, in turn, to other formulations of the invariant, including the original one by Fu, Kane and Mele [Cornean_2017].
4.1. Review: time-reversal symmetric matching matrices